OpenStudy (albert0898):

At the end of a meeting all participants shook hands with each other. Twenty-eight handshakes were exchanged. How many people were at the meeting?

OpenStudy (hossamhoussien):

@Albert0898 Assume that there were x people there and there were y handskakes and everybody in the meeting handshakes with everyone BUT him so \[x(x-1)=y\] but this count that everyone shakes twice \[\frac{x(x-1)}{2}=y\] \[x^2 - x-56=0\] \[x=8 ,-7 \] but (-7) is rejected so the answer will be 8 people